OpenStudy (skyval):

A ladder is placed 30 inches from a wall. It touches the wall at a height of 50 inches from the ground. The angle made by the ladder with the ground is ______ degrees, and the length of the ladder is _____ inches.

OpenStudy (matlee):



OpenStudy (matlee):

o gosh i dont know this who wanna take over lol

OpenStudy (matlee):

wait nvm i think i got it

OpenStudy (matlee):

Ok so u wanna find the legnth and angle of the ladder , you know how you place a ladder down its away from the way and makes it way to the wall

OpenStudy (matlee):

Length, would be found by using the pythogoreum theroem which is not so hard a^2 +b^2=c^2

OpenStudy (matlee):

^2 = squared like when you multiply 8 squared its 64

OpenStudy (matlee):

So in this case u have to use it in this equation a^2+b^2=c^2

OpenStudy (matlee):

u still here

OpenStudy (matlee):

Ok yay

OpenStudy (skyval):

Yes, I understand the pythagorean theorem

OpenStudy (matlee):

So in this case how would you put the Pythagorean theorem into this equation. For example find A, B,C

OpenStudy (skyval):

Okay i understand, so just plug in right

OpenStudy (matlee):

Yes exactly, but not for the angle just the length. Do you know which is a b and c

OpenStudy (matlee):

?